A researcher claims that adult hogs fed a special diet will have an average weight of 200 pounds. A sample of 10 hogs has an average weight 198.2 pounds and standard deviation of 3.3 pounds. At a= 0.05, can the claim be rejected? Also, find the 95% confidence interval of the true mean?

1
Expert's answer
2022-05-12T14:44:18-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 9" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.262156."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(198.2-2.262156\times\dfrac{3.3}{\sqrt{10}},"

"198.2+2.262156\times\dfrac{3.3}{\sqrt{10}})"

"=(195.839,200.561)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "195.839 < \mu < 200.561

," which indicates that we are 95% confident that the true population mean "\mu"  is contained by the interval "(195.839, 200.561)."

The following null and alternative hypotheses need to be tested:

"H_0:\mu=200"

"H_a:\mu\not=200"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=9" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.262156."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.262156\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{198.2-200}{3.3/\sqrt{10}}=-1.7249"

Since it is observed that "|t| = 1.7249 \le 2.262156=t_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed "df=9" degrees of freedom, "t=-1.7249" is "p=0.118631," and since "p=0.118631>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 200, at the "\alpha = 0.05" significance level.